Write a system of equations and solve:
The line with equation y+ax=c, passes through the points (1,5) and (3,1). Find a and c.

Answer by lwsshak3(11628)

Write a system of equations and solve:
The line with equation y+ax=c, passes through the points (1,5) and (3,1). Find a and c.
***
equation of a line: y=mx+b, m=slope, b=y-intercept
y=-ax+c
slope,a=∆y/∆x=(1-5)/(3-1)=-4/2=-2
equation:y=-2x+c
solve for c using coordinates of one of given points (1,5)
5=-2+c
c=7
equation: y=-2x+7